# RAFT: Robust Augmentation of FeaTures for Image Segmentation

### v1

@@ 2.1.2 Data Augmentation Methods
In addition to algorithmic approaches, data augmentation is another technique for dealing with data imbalance. Traditional augmentation methods, such as geometric transformations (rotation, flipping, cropping, and scaling) and photometric adjustments (brightness, contrast, and color alterations), have been widely used to improve model generalization by increasing dataset diversity [2]. Deep learning-based augmentation methods leverage generative models [15, 14, 34] to synthesize new data points that capture complex variations beyond simple transformations. These generative models have been successfully applied in medical imaging [26, 4, 9, 42] and underwater object recognition[23, 24], demonstrating their potential to generate realistic synthetic samples that enhance model robustness.
Recent advancements in augmentation strategies extend beyond raw image transformations to feature space modifications. Methods such as feature-based augmentation [20] [39] introduce diversity at a more abstract level, leveraging learned feature embeddings to generate novel training samples. This approach has been particularly effective in semi-supervised learning settings, where labeled data is scarce [19].
Hyperbolic neural networks often struggle to generalize when trained on few-shot, limited datasets. HFA addresses this issue by leveraging feature augmentations in hyperbolic space. Specifically, HFA generates class-identity-preserving features by modeling their distribution with a per-class wrapped normal distribution on the hyperbolic manifold. To accurately estimate the parameters of each distribution—including hyperbolic curvature, mean, and covariance—HFA employs a meta-learning framework based on neural ordinary differential equations (ODEs). In this framework, the iterative update of distribution parameters is modeled as a continuous gradient flow, which is then solved via the RK4 [3] method. This neural ODE-based gradient flow network leverages prior knowledge to achieve a more precise approximation of the underlying distribution even in data-scarce regimes.
@@ 3.2.1 Hyperbolic Mixup
We combine these mixed embeddings with the sampled embeddings from our learned class distributions into a single augmentation pool . When reintegrating these features into the spatial feature map, we randomly select either the mixed or sampled embeddings on a per-class basis.
@@ 5.1 Comparison With the State-of-the-Art
Examining the segmentation masks generated by RAFT and HALO using SegFormer B4 on the Cityscapes validation split, shown in Figure 2, while both segmentations are generally high quality, one notices that where HALO appears to struggle with the hood and its ornament of the car the photos are being captured from, RAFT has comparatively little trouble in ignoring it, with the obvious exception of the first photo, in which it misclassifies the hood ornament as being a bicycle, and a small part of the hood as being a person. Additionally, within the second photo, RAFT misclassifies a small sliver of the hood as being sky pixels. A smaller but still noticeable area where RAFT improves over HALO is in the sidewalk the pedestrians are walking to in photo 2. While the RAFT-trained SegFormer B4 still doesn’t fullly classify the sidewalk correctly, it classifies more of the overall shape compared to the HALO-trained SegFormer B4.

### v2

@@ 2.1.2 Data Augmentation Methods
Data augmentation is a key strategy for addressing data imbalance and improving model generalization. Traditional techniques like geometric and photometric transformations increase dataset diversity [2], while deep learning-based methods use generative models [15, 14, 34] to synthesize realistic samples, proving effective in domains like medical imaging [26, 4, 9, 43] and underwater object recognition [23, 24]. Recent approaches augment data in feature space [20, 40], generating novel training samples from learned embeddings—particularly beneficial in semi-supervised learning where labeled data is limited [19].
Hyperbolic neural networks often struggle with few-shot generalization. HypMix [37] addresses this by extending mixup into hyperbolic space: inputs are mapped to hyperbolic space, interpolated, and then mapped back to Euclidean space. To handle unlabeled data, the authors propose Möbius Gyromidpoint Label Estimation (MGLE), which generates augmented inputs, maps their logits into hyperbolic space, interpolates them, and converts the result back to Euclidean space to form pseudolabels. These are then combined with labeled data for HypMix training.
Another method to tackle this issue of limited data is HFA. Specifically, HFA generates class-identity-preserving features by modeling their distribution with a per-class wrapped normal distribution on the hyperbolic manifold. To accurately estimate the parameters of each distribution—including hyperbolic curvature, mean, and covariance—HFA employs a meta-learning framework based on neural ordinary differential equations (ODEs). In this framework, the iterative update of distribution parameters is modeled as a continuous gradient flow, which is then solved via the RK4 [3] method. This neural ODE-based gradient flow network leverages prior knowledge to achieve a more precise approximation of the underlying distribution even in data-scarce regimes.
@@ 3.2.1 Hyperbolic Mixup
We combine these mixed embeddings with the sampled embeddings from our learned class distributions into a single augmentation pool . When reintegrating these features into the spatial feature map, we randomly select either the mixed or sampled embeddings on a per-class basis. Unlike HypMix [37], we not only perform augmentation at the pixel-embedding level of granularity, but we also incorporate synthetic, generated augmentation directions, in order to create more diverse features.
@@ 5.1 Comparison With the State-of-the-Art
Examining the segmentation masks generated by RAFT and HALO using SegFormer B4 on the Cityscapes validation split, shown in Figure 2, while both segmentations are generally high quality, one notices that where HALO appears to struggle with the hood and its ornament of the car the photos are being captured from, RAFT has comparatively little trouble in ignoring it, with the obvious exception of the first photo, in which it misclassifies the hood ornament as being a bicycle, and a small part of the hood as being a person. Additionally, within the second photo, RAFT misclassifies a small sliver of the hood as being sky pixels. A smaller but still noticeable area where RAFT improves over HALO is in the sidewalk the pedestrians are walking to in photo 2. While the RAFT-trained SegFormer B4 still does not fully classify the sidewalk correctly, it classifies more of the overall shape compared to the HALO-trained SegFormer B4.